OpenStudy (anonymous):

Write the Maclaurin series for f(x)= e^x. I do not understand why its first term is "1".

OpenStudy (anonymous):

f(x) = e^x f' (x) = e^x f^n (x) = e^x no matter how often you differentiate it, it is still e^x and the formula an = f^n (x) / n/ = where x equals zero f^n (0 ) / n/ = e^0 / n! = 1 /n! and a1 = f (0 ) = 1 Maclaurin series of f(x) = e^x = 1 + x + x²/2/ + x³/3/ + ... ( / means factorial)

OpenStudy (anonymous):

I was confused as to why 0! =1. I should remember it like an exponent: 8^0 = 1. I understand now, thanks

OpenStudy (zarkon):

0! by definition is 1

OpenStudy (anonymous):

yea i totally understand your confusion. 0! is 1 , anything to the power of 0 is also 1

OpenStudy (anonymous):

Thanks guys! There may be more questions about Maclaurin